OpenStudy (anonymous):

x minus 5 all over 4 = 3.

OpenStudy (unklerhaukus):

\[\frac{x-5}4=3\] multiply both sides of the equation by 4 then add 5 to both sides

OpenStudy (anonymous):

(x - 5)/4 = 3 x - 5 = 12 Calculate x from this...

OpenStudy (anonymous):

Um is it 17?

OpenStudy (anonymous):

Yup :)